OpenStudy (bizzeizzy):

Simplify the following expression: |2+4i|

OpenStudy (bizzeizzy):

\[6i\] \[-6i\] \[\sqrt{8}\] \[\sqrt{20}\]

OpenStudy (freckles):

\[|a+bi|=\sqrt{a^2+b^2}\]

OpenStudy (michele_laino):

and:\[a=2, b=4\]

OpenStudy (michele_laino):

please note that: \[i=\sqrt{-1}\] @familyguymath

OpenStudy (anonymous):

opps